OpenStudy (anonymous):

tan^-1[(√x+√a)/(1-√(xa))] differentiate it plz m gtting 1/2(√x) which is not correct answer is 1/[2(√x)(1+x)]

hartnn (hartnn):

same type of question with different substitution

OpenStudy (anonymous):

hmm i think i did something wrong let me do it again

OpenStudy (anonymous):

well u don't have to substitue :) @hartnn i think i forgot to differentiate tan^-1 (x)^1/2 i missed the tan^-1 part

hartnn (hartnn):

say, sqrt x= tan y sqrt a = tan alpha then , [(√x+√a)/(1-√(xa))] =... ?

hartnn (hartnn):

ohh..so now did u get it ?

OpenStudy (anonymous):

u have to use tan^-1 x + tan^-1 y formula well not now

OpenStudy (anonymous):

trying again

OpenStudy (anonymous):

:'( not getting :'(

OpenStudy (anonymous):

how to differnatiate tan^-1(x)^1/2

hartnn (hartnn):

use chain rule, know about it ?

OpenStudy (anonymous):

no :'( that's the problem simple chain rule i know

OpenStudy (anonymous):

like this one sin(cos3x) but never did for inverse trign.........

OpenStudy (anonymous):

is the differenatition of tan^-1(x)^1/2 is

OpenStudy (anonymous):

is =X =1/(1+(x^1/2)^2 . 1/2*(x)^-1/2 ????????

hartnn (hartnn):

ok, derivative of tan^{-1} x is 1/(1+x^2), right ?? for derivative of tan^{-1} sqrt(x) you do 1/(1+ [\sqrt x]^2) times deribative of sqrt x

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

i think i have used chain rule -_- is that a chain rule @hartnn

OpenStudy (souvik):

you did it,,,kitty

hartnn (hartnn):

didn't you get the answer then ? yes, thats was chain rule only...

OpenStudy (anonymous):

lol i know everything but still do silly things -_- @souvik

OpenStudy (souvik):

it happens some time...

OpenStudy (anonymous):

@hartnn - chain rule :O so that's a chain rule

hartnn (hartnn):

because what you wrote matches with your final answer...

OpenStudy (anonymous):

not some times everytime it happens

OpenStudy (souvik):

keep your head cool....like MSD

OpenStudy (anonymous):

bcoz of calculation mistake i lost 3 marks in maths prob :'( i did everything correct but in final answer i made a error in multiplication

OpenStudy (anonymous):

MSD??????

OpenStudy (anonymous):

@hartnn -tahnx :)

hartnn (hartnn):

hey, welcome anytime ^_^

OpenStudy (anonymous):

:)

OpenStudy (souvik):

Indian cricket team captain...

OpenStudy (anonymous):

oh :OOO ms dhoni :OOO i don't like him lol

OpenStudy (souvik):

why?

OpenStudy (anonymous):

reminds me of my frenemy :P